OpenStudy (anonymous):

I need help to Evaluate the following logarithm: log[(1)/(3)](0.384). Rewrote in equation form in the comments!

OpenStudy (anonymous):

\[\log_{1/3} (0.384)\]

OpenStudy (amistre64):

\[\frac{L(384/1000)}{L(1/3)}\] \[\frac{L(384)-L(1000)}{L(1)-L(3)}\] \[-\frac{L(384)-L(10^3)}{L(3)}\] \[-\frac{L(384)-3L(10)}{L(3)}\]

OpenStudy (amistre64):

what constitutes an evaluation?

OpenStudy (anonymous):

The question asked me to evaluate the logarithm expression and round the answer to two decimal places.